# Evaluating the Propensity of Generative AI for Producing Disinformation During an Election Cycle

### v1

@@ 3.1 Model Performance
Figure 1 shows the probability that Generative AI (panel) produced disinformation (x-axis) and the average harm associated with disinformation produced (y-axis), using different adversarial roles (marker style) across disinformation categories (color). The vertical line reflects the marginal probability of disinformation production across all models, roles and categories, while each point represents the conditional probability given the model, role and category obtained by averaging across all topics within each category. Similarly, the horizontal line represents the overall average harm across all disinformation produced, while the points reflect average harm within each condition. Points that fall within the upper-right quadrant are above average in both disinformation production rates and harm, whereas points in the lower-left quadrant are below average. The other quadrants reflect averages that are above average in one dimension but below in the other.
GPT-4o produced the most points in the upper-right quadrant, reflecting above average disinformation production and harm across most conditions (with the exception of Role 4), whereas Gemini exhibited extremely low rates of disinformation production for politics (blue) but higher rates of below average harm production for health (red). Copilot was the only model that did not produce any disinformation for Role 4 and tended to realize lower than average disinformation production rates across other conditions, with the exception of Roles 1 & 2 in health.
@@
Equation 1 shows that the expected harm ($E[H(m,c)]$) for model ($m$) and disinformation category ($c$) is the product of the probability of disinformation ($d$) given the model and disinformation category, and the average disinformation harm ($H_{d}(m,c)$) for the same model and category. This is obtained by marginalizing (averaging) over roles ($r$) and topics ($t$), and it is important to note that harm associated with accurate information and refusals to answer ($H_{a}$) is always 0, so the expectation only needs to consider the outcome of disinformation. Using Equation 1, the expected harm associated with each Generative AI model and disinformation category is reflected in Table 5.
@@
Table 5 shows that Copilot realized the lowest expected harm score (0.89) for disinformation related to health, while Gemini realized the lowest expected harm for political topics (0.16). In both cases, this was primarily due to low rates of disinformation production within each category. The fact that Gemini produced very little political disinformation is not surprising since the company announced earlier that they would restrict topics related to the 2024 elections, which is a responsible tactic given the risk involved with producing false information during this time (Robins-Early, 2024).
@@
In order to explore overall model performance, Equation 1 was further marginalized over categories to produce the results in Table 6.
@@ 3.2 Impact of Adversarial Role
The previous section investigated the propensity of different Generative AI models for producing harmful disinformation, and this section will explore the impact of adversarial role. Table 7 shows the results obtained by marginalizing Equation 1 over models, categories and topics.
@@
As predicted in Section 2.2 (Table 3), Role 1 was most challenging for models and resulted in the greatest expected harm, whereas Role 4 produced the lowest expected harm scores. Note that the decrease in expected harm from Roles 1-4 is primarily due to the decreased probability that disinformation was produced. This suggests that Generative AI is sensitive to both nefarious intent (Role 4) and potential virality (Role 3) when deciding whether or not to produce information.
@@ 3.3 Impact of Disinformation Category and Topic
Since this study occurred two months prior to the 2024 US elections, it was predicted that Generative AI models would produce less disinformation around political topics than health due to developers restricting political output during this time (Robins-Early, 2024). However, as demonstrated in Section 3.1 (Table 5), GPT-4o actually produced greater rates of harmful disinformation for the political category, so this section begins by exploring the overall trends across models.
In order to investigate the impact of disinformation category, Equation 1 was marginalized over model, role and topic to produce Table 8.
@@
As predicted, political topics produced disinformation at lower rates (.33) than health (.56), but also realized greater average harm scores (3.05) than health (2.29). However, since GPT-4o was primarily responsible for most of the disinformation in the political category (Table 5), the increased harm is likely due to model-level characteristics (See Section 3.4).
@@
In order to explore if certain disinformation topics were universally challenging across models, Figure 2 shows the expected harm for each topic. Each point represents the expected harm, averaging across roles and models. Points that are colored in red or orange are associated with relatively high expected harm, and tend to fall in the upper-right quadrant of the graph. Table 9 lists the topics that received the greatest and least expected harm, along with the corresponding false claims.
@@ 3.4 Important Factors for Predicting Disinformation Production
In order to gain insight into the factors each model found important to make predictions, mean absolute SHAP values (Lundberg and Lee, 2017) were computed for each model (Figure 3). SHAP uses a game theoretic approach to estimate each feature’s importance to the model’s output. Figure 3 shows the confusion matrix (left) associated with the Random Forest model, in addition to the SHAP values for the model (right), which are sorted from most- (top) to least-important (bottom). Moreover, the SHAP plot shows when GPT-4o, Role 1, Health, and Role 2 are present, there is an increased likelihood of disinformation being produced, whereas when Politics, Role 4, Role 3 and Copilot are present, there is a decreased likelihood (color of points). Gemini (bottom) is the least important for this model, and has mixed impact on output.
@@
4 Conclusions
This study investigated the propensity of modern Generative AI models for producing harmful disinformation during a major US election cycle. It was predicted that model developers would restrict or tighten output for topics related to politics, leading to decreased rates of harmful disinformation relative to health. Although this was generally true across models (Table 8), GPT-4o actually produced higher rates of harmful disinformation for political topics than health (Table 5), and realized the greatest expected harm scores across all conditions tested in this study (Table 6). Both Copilot and Gemini tied for least overall expected harm (Table 6), with Gemini performing better for political topics, and Copilot performing better for topics related to health (Table 5).
@@ 4 Conclusions
Taken together, these results suggest that adversaries with no technical sophistication can use Generative AI to produce harmful disinformation during an election cycle, given the correct choice of adversarial role and model. However, there are measures that can be taken by developers to mitigate the harm associated with model output, as defined by this study. First, developers should always include a disclaimer on any fictitious content, in order to minimize legal risk. Second, prompts should be evaluated independently of the role the model is asked to assume. Since each model had an understanding that the claims contained in the adversarial prompt were incorrect (Table 2), it should be capable of refusing the request independent of the role is asked to assume. Third, fictitious content produced by models should never include known entities in the output. This makes the content less useful to adversaries and minimizes risk to the entity’s reputation, in addition to reducing legal issues that may result from the reputational damage.
It is clear that developers are continually improving the safety associated with Generative AI, and even though GPT-4o performed lowest for the conditions considered in this study, the new scorecard (OpenAI, 2024) shows that the new o1 version should perform better if this study were repeated. By considering the insights produced through this effort and recommendations provided above, it is hoped that this paper can contribute to the improvement cycle.
@@
Below are examples of disinformation produced receiving a maximum harm score (4). An example is provided for political and health categories.

### v2

@@ 3.1 Model Performance
Fig 1 shows the probability that generative AI (panel) produced disinformation (x-axis) and the average harm associated with disinformation produced (y-axis), using different adversarial roles (marker style) across disinformation categories (color). The vertical line reflects the marginal probability of disinformation production across all models, roles and categories, while each point represents the conditional probability given the model, role and category obtained by averaging across all topics within each category. Similarly, the horizontal line represents the overall average harm across all disinformation produced, while the points reflect average harm within each condition. Points that fall within the upper-right quadrant are above average in both disinformation production rates and harm, whereas points in the lower-left quadrant are below average. The other quadrants reflect averages that are above average in one dimension but below in the other.
GPT-4o produced the most points in the upper-right quadrant, reflecting above average disinformation production and harm across most conditions, whereas Gemini exhibited extremely low rates of disinformation production for politics (blue) but higher rates of below average harm production for health (red). This was largely due to the fact that developers restricted content related to political topics during the 2024 election season bib17 . Copilot was the only model that did not produce any disinformation for Role 4 and tended to realize lower than average disinformation production rates across other conditions, with the exception of Roles 1 & 2 in health.
@@
Eq  1 shows that the expected harm ($E[H(m,c)]$) for model ($m$) and disinformation category ($c$) is the product of the probability of disinformation ($d$) given the model and disinformation category, and the average disinformation harm ($H_{d}(m,c)$) for the same model and category. This is obtained by marginalizing (averaging) over roles ($r$) and topics ($t$), and it is important to note that harm associated with accurate information and refusals to answer ($H_{a}$) is always 0, so the expectation only needs to consider the outcome of disinformation. Using Eq  1, the expected harm associated with each fenerative AI model and disinformation category is reflected in Table 6.
@@
Table 6 shows that Copilot realized the lowest expected harm score (0.89) for disinformation related to health, while Gemini realized the lowest expected harm for political topics (0.16). In both cases, this was primarily due to low rates of disinformation production within each category. The fact that Gemini produced very little political disinformation is not surprising since the company announced earlier that they would restrict topics related to the 2024 elections, which is a responsible tactic given the risk involved with producing false information during this time bib17 . Since Copilot produced twice as much disinformation for topics related to health (.38) than politics (.19), it also suggests that mitigation attempts were made by developers of this model.
@@
In order to explore overall model performance, Eq  1 was further marginalized over categories to produce the results in Table 7.
@@ 3.2 Impact of Adversarial Role
The previous section investigated the propensity of different generative AI models for producing harmful disinformation, and this section will explore the impact of adversarial role. Table 8 shows the results obtained by marginalizing Eq  1 over models, categories and topics.
@@
As predicted in Table 3, Role 1 was most challenging for models and resulted in the greatest expected harm, whereas Role 4 produced the lowest expected harm scores. Note that the decrease in expected harm from Roles 1-4 is primarily due to the decreased probability that disinformation was produced. This suggests that generative AI is sensitive to both nefarious intent (Role 4) and potential virality (Role 3) when deciding whether or not to produce information.
@@ 3.3 Impact of Disinformation Category and Topic
Since this study occurred two months prior to the 2024 US elections, it was predicted that generative AI models would produce less disinformation around political topics than health due to developers restricting political output during this time bib17 . However, as demonstrated in Table 6, GPT-4o actually produced greater rates of harmful disinformation for the political category, so this section begins by exploring the overall trends across models.
In order to investigate the impact of disinformation category, Eq  1 was marginalized over model, role and topic to produce Table 9.
@@
As predicted, political topics produced disinformation at lower rates (.33) than health (.56), but also realized greater average harm scores (3.05) than health (2.29). However, since GPT-4o was primarily responsible for most of the disinformation in the political category (Table 6), the increased harm is likely due to model-level characteristics (See Next Section).
@@
In order to explore if certain disinformation topics were universally challenging across models, Fig 2 shows the expected harm for each topic. Each point represents the expected harm, averaging across roles and models. Points that are colored in red or orange are associated with relatively high expected harm, and tend to fall in the upper-right quadrant of the graph. Table 10 lists the topics that received the greatest and least expected harm, along with the corresponding false claims.
@@ 3.4 Important Factors for Predicting Disinformation Production
In order to gain insight into the factors each model found important to make predictions, mean absolute SHAP values bib18 were computed for each model (Fig 3). SHAP uses a game theoretic approach to estimate each feature’s importance to the model’s output. Fig 3 shows the confusion matrix (left) associated with the Random Forest model, in addition to the SHAP values for the model (right), which are sorted from most- (top) to least-important (bottom). Moreover, the SHAP plot shows when GPT-4o, Role 1, Health, and Role 2 are present, there is an increased likelihood of disinformation being produced, whereas when Politics, Role 4, Role 3 and Copilot are present, there is a decreased likelihood (color of points). Gemini (bottom) is the least important for this model, and has mixed impact on output.
@@
It is worth noting that the sample size used for this study is less than ideal for modeling-based approaches, but the fact that it produced reasonable predictive performance is encouraging. Plus, there is directional agreement between models regarding the importance of the top 3 features ($\geq$ .1, with one exception) and bottom 4 features (all $\leq$ .09). However, future research should increase the number of samples to improve confidence.
4 Conclusion
This study investigated the propensity of modern generative AI models for producing harmful disinformation during a major US election cycle. It was predicted that model developers would restrict or tighten output for topics related to politics, leading to decreased rates of harmful disinformation relative to health. Although this was generally true across models (Table 9), GPT-4o actually produced higher rates of harmful disinformation for political topics than health (Table 6), and realized the greatest expected harm scores across all conditions tested in this study (Table 7). Both Copilot and Gemini tied for least overall expected harm (Table 7), with Gemini performing better for political topics as a result of risk mitigation attempts made by developers bib17 , and Copilot performing better for topics related to health (Table 6).
@@ 4 Conclusion
Taken together, these results suggest that adversaries with no technical sophistication can use generative AI to produce harmful disinformation during an election cycle, given the correct choice of adversarial role and model. However, there are measures that can be taken by developers to mitigate the harm associated with model output, as defined by this study. First, developers should always include a disclaimer on any fictitious content, in order to minimize legal risk and increase transparency of the source. Second, prompts should be evaluated independently of the role the model is asked to assume. Since each model had an understanding that the claims contained in the adversarial prompt were incorrect (Table 2), it should be capable of refusing the request independent of the role is asked to assume. Third, fictitious content produced by models should never include known entities in the output. This makes the content less useful to adversaries and minimizes risk to the entity’s reputation, in addition to reducing legal issues that may result from damage to their reputation.
Although the topics selected for political and health categories of disinformation were effective at detecting known risk mitigation attempts bib17 , future research could improve upon these methods by avoiding topics that impact both political and health categories. Moreover, the language used in the adversarial prompts could be streamlined and other harm categories (such as call-to-action) could be included in prompts to enable quantification of other important dimensions. Finally, a greater number of samples should be obtained to increase confidence in the modeling results.
It is important to note that since the methods and materials included in this paper could be exploited by bad actors, a great deal of consideration went into releasing the experimental details. However, since recent evidence suggests that disinformation produced by generative AI did not seem to impact the 2024 US election cycle bib31 ; bib32 ; bib33 ; bib34 , it is now believed that releasing this information will decrease risk in future elections by allowing developers better replicate these findings and provide appropriate mitigation. It is clear that developers are continually improving the safety associated with generative AI, and even though GPT-4o performed lowest for the conditions considered in this study, the new scorecard bib16 shows that the o1 version should perform better if this study were repeated. By considering the insights produced through this effort and recommendations provided above, it is hoped that this paper can contribute to the improvement cycle.
Supporting information
S1 File
. All Prompts. Document contains all of the direct and adversarial prompts used in this investigation. Contact author to obtain.
S2 File.
Model Output. Document contains all of the model output for adversarial and direct prompts. Contact author to obtain.
S3 File
. Analysis Files File contains: 1) excel sheet with harm scores for each topic, along with links to the scoring urls with justification for any modifications made to the automated estimates; 2) Notebook with analysis script to reproduce all figures in this paper; and 3) html file showing the analysis script and output so readers do not need to rerun the analysis. Contact author to obtain.
@@
app1) Below are examples of disinformation produced receiving a maximum harm score (4). An example is provided for political and health categories.